# Attosecond X-ray Chronoscopy of Core-level Photoemission

### v3

@@
Attosecond X-ray Chronoscopy of Core-level Photoemission
@@
Attosecond photoemission or photoionization delays are a unique probe of the structure and the electronic dynamics of matter. However, spectral congestion and spatial delocalization of valence electron wave functions set fundamental limits to the complexity of systems that can be studied and the information that can be retrieved, respectively. Using attosecond X-ray pulses from LCLS, we demonstrate the key advantages of measuring core-level delays: the photoelectron spectra remain atom-like, the measurements become element specific and the observed scattering dynamics originate from a point-like source. We exploit these unique features to reveal the effects of electronegativity and symmetry on attosecond scattering dynamics by measuring the photoionization delays between N-1s and C-1s core shells of a series of aromatic azabenzene molecules. Remarkably, the delays systematically increase with the number of nitrogen atoms in the molecule and reveal multiple resonances. We identify two previously unknown mechanisms regulating the associated attosecond dynamics, namely the enhanced confinement of the trapped wavefunction with increasing electronegativity of the atoms and the decrease of the coupling strength among the photoemitted partial waves with increasing symmetry. This study demonstrates the unique opportunities opened by measurements of core-level photoionization delays for unravelling attosecond electron dynamics in complex matter.
@@
Tracking attosecond multi-electron dynamics with atomically resolved spatial information is the current focus of numerous research efforts worldwide. X-ray spectroscopy has established itself as a powerful solution to breaking the complexity barriers faced by valence-shell spectroscopies. These barriers are particularly noticeable in time-resolved studies of the photoelectric effect, known as attosecond chronoscopy (?, ?, ?). This technique has integrated time-domain access to multi-electron dynamics into the powerful framework of electron-scattering physics which has created a flourishing research field with applications to molecules (?, ?, ?, ?, ?, ?), clusters (?, ?), liquids (?) and solids (?, ?, ?, ?). Such studies have traditionally made use of attosecond pulses in the extreme ultraviolet (up to 120 eV), which has so far prevented access to atomic core levels.
@@
Here, we demonstrate the unique opportunities that arise from advancing the research field of attosecond chronoscopy from the extreme-ultraviolet to the X-ray domain. This adds element specificity and site selectivity to the method, while simultaneously confining the emission site of the photoelectron wave to a point-like source. Moreover, core-shell photoelectron spectra have a much simpler structure than valence-shell spectra because they feature a single band per atom of each type (when neglecting satellite lines), which facilitates the extension of the method to complex forms of matter. This advance was made possible by groundbreaking progress in free-electron-laser (FEL) science that has led to the generation of attosecond pulses in the X-ray domain through the XLEAP technique (?). The long-standing challenge of temporal jitter between the FEL pulses and laser pulses was solved by using attosecond angular streaking with a circularly-polarized infrared laser pulse and detecting the photoelectron momentum distribution in the polarization plane (?, ?).
@@
The present experiments establish the unique capabilities of X-ray chronoscopy at unraveling the mechanisms of attosecond electron dynamics in complex systems, notably in aromatic molecules (i.e., benzene derivatives). Specifically, our study reveals the multiple-scattering phenomena that lead to transient trapping of photoelectrons following their release from specific atoms. Our results reveal two previously unknown mechanisms regulating such attosecond scattering dynamics. First, we find that substituting the ionized functional group with a more electronegative one (C-H N) increases the trapping times. Second, we show that increasing the symmetry of the molecular scaffold also increases the trapping times by suppressing the coupling to the less-trapped partial waves with lower angular momentum. To demonstrate these effects, we chose the azabenzene molecules pyridine, pyrimidine and -triazine, depicted in Fig. 1A. We produced attosecond X-ray pulses with energies just above the nitrogen K-edge located at 405 eV and measured the photoionization delays between the electrons emitted from the nitrogen -shell (N-1s) and those emitted from the carbon -shell (C-1s). Since the latter have kinetic energies of more than 110 eV, the C-1s electrons are emitted with negligible photoionization delays ( as), defining a self-referenced technique for accessing absolute photoionization delays of the N-1s electrons. We measured N-1s photoionization delays of up to 300 as close to the nitrogen -shell threshold, which generally decrease with increasing kinetic energy but display local maxima at 3-15 eV above the threshold. The measured delays also markedly increase from pyridine to pyrimidine and -triazine. These experimental results are compared with state-of-the-art core-level photoionization calculations, which reveal the presence of several shape resonances in this low kinetic-energy region. The comparison of experiment and theory reveals that the photoelectrons originating from the nitrogen -shell are trapped with increasing efficiency in the molecular plane in the sequence pyridine, pyrimidine, -triazine, i.e., with increasing number of nitrogen atoms. Additionally, the increasing symmetry of the molecular scaffold (from C in pyridine to D in -triazine) reduces the coupling among the photoelectron partial waves, which frustrates the decay of the trapped partial waves with high angular momentum. Pyrimidine is a building block of the nucleobases thymine, cytosine and uracil. All three azabenzene molecules are widespread structural motifs in biomolecules, drugs and molecular optoelectronics (?, ?, ?, ?). The discovered mechanisms therefore suggest guidelines for understanding electronic dynamics in biomolecules and molecular optoelectronics.
@@
Figure 1 illustrates the principle of the present experiments, carried out at the time-resolved atomic, molecular, and optical science (TMO) experimental hutch (?) of the Linac Coherent Light Source (LCLS). Linearly polarized attosecond X-ray pulses with a tunable central energy were obtained using the XLEAP technique (?) and were superimposed with circularly polarized short-wave infrared (IR, 1.3 m, 40 fs) pulses in the interaction region of a co-axial velocity-map-imaging (c-VMI) spectrometer. A continuous molecular beam containing pyridine, pyrimidine or -triazine intersected with the overlapping X-ray and IR pulses in the interaction region. The photoelectrons created from nitrogen -shell ionization of the sample were streaked by the IR pulses and projected onto a position-sensitive imaging detector with a central hole. A photoelectron image recorded by ionizing pyrimidine with X-ray pulses centered at 454 eV is shown in Fig. 1A, whereby the inner (outer) ring corresponds to ionization of the nitrogen (carbon) -shell electrons. Figures 1B, 1C, and 1D show how the relative photoionization delays between the N-1s and C-1s photoelectrons were determined. Because each photoelectron was angularly streaked into the direction of the IR-laser vector potential at the instant of photoionization (inset of Fig.1A), the relative streaking angle between N-1s and C-1s electrons directly encodes their relative photoionization delay. The conversion from relative streaking angle to the relative photoionization delay is simply given by , where is the period of the streaking laser (?, ?, ?, ?, ?, ?, ?).
@@
Figure 2 illustrates how the relative streaking angles were obtained from the experimental data. Due to the large temporal jitter (500 fs) between the arrival time of XFEL and IR laser pulses (?), the streaking directions of the core-shell photoelectrons were completely random from shot to shot. The information of relative photoionization time delay is therefore only encoded in the relative streaking angle of the N-1s and C-1s electrons, which is recorded on a shot-by-shot basis. Figure 2A shows the difference image between the average over all shots corresponding to a narrow range of streaking directions sorted by post-selection of the images and the average over all streaked images without any sorting and selection. An angular shift between the streaking directions of C-1s (outer) and N-1s (inner) electrons can be observed, which directly reflects the relative photoionization delays, as described above. To accurately determine the relative streaking angle, we employ a partial covariance analysis that utilizes the large shot-to-shot variation of streaking directions to directly extract without estimating and sorting the streaking directions of C-1s and N-1s electrons in every single image. For a given relative streaking angle , the streaking-induced signal fluctuations in C-1s electrons at have the most positive covariance with those in N-1s electrons around . By correctly calculating the partial covariance coefficients between streaking-induced signal changes in N-1s and C-1s electrons, the differential streaking angle manifests itself as a global shift in the partial covariance maps along the diagonal, as shown in Fig. 2C. Additional details on the partial covariance analysis are given in Section 2.2.1 of the supplementary material (SM). By comparing Figs. 3A,C to B,D, a relative photoionization delay of up to as is observed at eV, as opposed to a negligible delay at eV. We have also developed a complementary data-analysis method (the so-called ”center-of-mass” method), which is described in Section 2.2.2 of the SM. These two independent analysis methods yielded consistent photoionization delays (Fig. S14), which confirms the reliability and robustness of both methods.
@@
Since such shape resonances are notoriously sensitive to electron correlations and molecular structure (?, ?, ?), we have carefully benchmarked the core-level photoionization calculations against measured photoionization cross sections (Figs. 3D and S22), as well as photoelectron angular distributions measured in the present work (Figs. 3B and S23). Comparing the calculations using the equivalent-core (ECO) method with the experimental photoionization cross section (Fig. 3D) revealed very good agreement for the resonance lying below the nitrogen K-edge and the lowest-lying shape resonance after shifting the calculations to lower energies by 4.3 eV. In comparison, the correlation-polarization (CP) calculations only needed to be shifted by 0.6 eV to agree with the experimental cross sections and the calculated width of the lowest-lying shape resonance was in better agreement with the experiment than the ECO calculations. The CP calculations also agree better with the angular asymmetry parameters measured in this work (Figs. 3B and S23). Since both calculations overestimate the energetic position of the upper shape resonance by 3.5 eV, we therefore identified the CP calculations as the more accurate ones and from hereon exclusively discuss these results.
@@
Figures 3C,D demonstrate the two shape resonances of pyrimidine within the energy range, which lead to the local maxima of the cross section. The A symmetry component is dominating in the cross section, which displays an anti-bonding character and corresponds to the electron trapping that occurs in the molecular plane in both resonances. The insets in Fig. 3C show the one-electron representations of the corresponding resonant wavefunctions (Kapur-Peierls states, see SM). The lower-lying shape resonance features 5 nodal planes and the upper one 6, showing that partial waves of high angular momentum are dominating the photoelectron trapping in both cases. The narrower width of the lower resonance nicely correlates with a longer photoemission delay and the broader linewidth of the upper one correlates with the shorter delay. The partial cross section corresponding to a total final-state wavefunction of A symmetry, i.e., anti-symmetric with respect to the molecular plane and thus a -type continuum wavefunction, does not feature any sharp resonances, indicating that photoelectron trapping exclusively occurs in the molecular plane of pyrimidine. The probability density of the resonant wavefunction of -triazine at 418 eV is illustrated in the inset of Fig. 3E and in Fig. 4. Similar resonances occur in all three molecules, as shown in Fig. S20.
@@
Since our measurements extend to within 2 eV of the N-1s ionization threshold, they are particularly sensitive to the CLC delays. It has previously been shown that CLC delays in attosecond streaking experiments are equivalent to the continuum-continuum (cc) delays known from RABBIT measurements (?). The most commonly used approximation, known as “P”, only includes the phase of the continuum-continuum (cc) matrix elements (?). The “P+A” formula additionally contains a long-range amplitude correction. Through detailed comparisons of these (and two more) formulae with calculations based on the time-dependent Schrödinger equation (TDSE, see Fig. S24A), we found that the arithmetic average of “P” and “P+A” agreed best with the TDSE. We therefore used those results in Figs. 3 and 4 of the present article and show the other results in the remainder of Fig. S24. We also investigated the role of the Auger decay and the post-collision interaction (PCI) between the Auger electron and the photoelectron. We find that the correction of the Coulomb delay from a singly- to a doubly-charged cation is nearly compensated by the delay that is caused by the PCI effect (Fig. S25). Since these corrections are small and do not affect the relative delays between the molecular species that we studied, we refer the interested reader to the SM, and neglect these corrections for the remainder of the present discussion.
@@
These results open the novel opportunity to analyze the mechanisms that govern the trapping of the photoelectron on the attosecond time scale. Figure 4 illustrates the trapping of the outgoing photoelectron wave in the higher-lying (418 eV) resonance of -triazine. The atomic cores provide attractive potentials in their vicinities, while the centrifugal energy creates a repulsive region at the center of the aromatic ring. Their addition results in a potential barrier (black curve in Fig. 4) which traps the resonant wavefunction for a duration of 360 as. The photoelectron wave initially originates from the 1s shells of the nitrogen atoms and then scatters in the molecular potential, which results in comparable amplitudes of the resonant wave function on all atoms of the molecular ring. But what causes the increase of the photoionization delays and resonance lifetimes with increasing number of nitrogen atoms and increasing symmetry?
@@
To answer this question, Figure 5A shows a sketch of effective radial potentials which indicate the height of the angular-momentum barrier for different values of the angular momentum, , and Figs. 5C,E,G show the absolute squares of resonant one-electron wavefunctions, resolved into partial waves of angular momentum and irreducible representations of the respective point groups (specified in each panel). Figure 5B shows the CLC delays and Figs. 5D,F,H show the calculated delays for all three molecules. Because of its isolated nature and visibility in the experimental results (Fig. 3C), we focus the remainder of this discussion on the shape resonance located at 417-420 eV in each molecule.
@@
In all three molecules, the partial wave(s) with dominate(s) in amplitude, which is the consequence of the high barrier in the radial potential (panel A) arising from the centrifugal-potential contribution , where is the mass of the electron and its distance from the center of charge. The effective potentials of are indicated by the shaded areas in the corresponding panels. It is clear that the main part of the wavefunction is confined in the quasi-bound region in the vicinity of the N and C atoms, while a tail tunnels through the barrier. In the case of pyridine (Fig. 5A), the dominant partial wave is least-well trapped compared to the other molecules, which manifests itself in the ratio of the absolute-square amplitudes of the resonant wavefunction between the local maximum at small and the local minimum indicated by the second grey arrow in Figs. 5C,E,G. This identifies tunneling through an angular-momentum barrier as the first mechanism governing electron trapping in the azabenzene series and shows that the trapping strength of the dominant partial wave increases in the order pyridine, pyrimidine, -triazine. This shows that substituting a C-H group with the more electronegative N atom increases the trapping strength and therefore the lifetime of the resonance. We note that the divergence of the Kapur-Peierls state for large is a consequence of their mathematical nature, specifically their complex-valued momentum, see SM Section 3.2.
Figure 5 also reveals a second mechanism at play. In addition to the changing amplitude ratio between the wavefunctions in the quasi-bound and tunneling regions, we also observe the manifestation of molecular symmetry. Whereas pyridine displays sizable contributions from an partial wave, which can be identified as being weakly trapped from the very low amplitude of its quasi-bound part, the contributions of are much weaker in pyrimidine and barely visible in -triazine as a consequence of symmetry. The three-fold (D) symmetry of the -triazine scaffold is indeed the reason why and partial waves are dominant and other partial waves are suppressed in that case. The suppression of these lower-, more weakly trapped partial waves, to which the partial waves can couple to release the photoelectron more efficiently, explains the increasing lifetime of the shape resonance in the series pyridine, pyrimidine, -triazine.

### v4

# Attosecond X-ray Core-level Chronoscopy of Aromatic Molecules

@@
Abstract
@@
Attosecond photoemission or photoionization delays are a unique probe of the structure and the electronic dynamics of matter. However, spectral congestion and spatial delocalization of valence electron wave functions set fundamental limits to the complexity of systems that can be studied and the information that can be retrieved, respectively. Using attosecond X-ray pulses from LCLS, we demonstrate the key advantages of measuring core-level delays: the photoelectron spectra remain atom-like, the measurements become element specific and the observed scattering dynamics originate from a point-like source. We exploit these unique features to reveal the effects of electronegativity and symmetry on attosecond scattering dynamics by measuring and calculating the photoionization delays between N-1s and C-1s core shells of a series of aromatic azabenzene molecules. Remarkably, the delays increase with the number of nitrogen atoms in the molecule and reveal multiple resonances. We identify two previously unknown mechanisms regulating the associated attosecond dynamics, namely the enhanced confinement of the trapped wavefunction with increasing electronegativity of the atoms and the decrease of the coupling strength among the photoemitted partial waves with increasing symmetry. This study demonstrates the unique opportunities opened by measurements of core-level photoionization delays for unraveling attosecond electron dynamics in complex matter.
I Introduction
@@
Tracking attosecond multi-electron dynamics with atomically resolved spatial information is the current focus of numerous research efforts worldwide. X-ray spectroscopy has established itself as a powerful solution to breaking the complexity barriers faced by valence-shell spectroscopies. These barriers are particularly pronounced in time-resolved studies of the photoelectric effect, known as attosecond chronoscopy [1, 2, 3]. This technique has integrated time-domain access to multi-electron dynamics into the powerful framework of electron-scattering physics which has created a flourishing research field with applications to molecules [4, 5, 6, 7, 8, 9], clusters [10, 11], liquids [12] and solids [13, 14, 15, 16]. Such studies have traditionally made use of attosecond pulses in the extreme ultraviolet (up to 120 eV), which has so far prevented access to atomic core levels.
@@
Here, we demonstrate the unique opportunities that arise from advancing the research field of attosecond chronoscopy from the extreme-ultraviolet to the X-ray domain. This adds element specificity and site selectivity to the method, while simultaneously confining the emission site of the photoelectron wave to a point-like source. Moreover, core-shell photoelectron spectra have a much simpler structure than valence-shell spectra, which facilitates the extension of the method to complex forms of matter. This advance was made possible by groundbreaking progress in free-electron-laser (FEL) science that has led to the generation of attosecond pulses in the X-ray domain through the X-ray laser-enhanced attosecond pulse generation (XLEAP) technique [17]. The long-standing challenge of temporal jitter between the FEL pulses and laser pulses was solved by using attosecond angular streaking with a circularly-polarized infrared laser pulse and detecting the photoelectron momentum distribution in the polarization plane [18, 19, 20].
@@
The present experiments establish the unique capabilities of X-ray chronoscopy, recently demonstrated in a diatomic molecule [20], at unraveling the mechanisms of attosecond electron dynamics in complex systems, notably in aromatic molecules (i.e., benzene derivatives). Specifically, our study reveals the multiple-scattering phenomena that lead to transient trapping of photoelectrons following their release from specific atoms. Our results reveal two previously unknown mechanisms regulating such attosecond scattering dynamics. First, we find that substituting the ionized functional group with a more electronegative one (C-H N) increases the trapping times. Second, we show that increasing the symmetry of the molecular scaffold also increases the trapping times by suppressing the coupling to the less-trapped partial waves with lower angular momentum. To demonstrate these effects, we chose the azabenzene molecules pyridine, pyrimidine and -triazine, depicted in Fig. 1A. We produced attosecond X-ray pulses with energies just above the nitrogen K-edge located at 405 eV and measured the photoionization delays between the electrons emitted from the nitrogen K-shell (N-1s) and those emitted from the carbon K-shell (C-1s). Since the latter have kinetic energies of more than 110 eV, the C-1s electrons are emitted with negligible photoionization delays ( as), defining a self-referenced technique for accessing absolute photoionization delays of the N-1s electrons. We measured N-1s photoionization delays of up to 300 as close to the nitrogen K-shell threshold, which generally decrease with increasing kinetic energy but display local maxima at 3-15 eV above the threshold. The measured delays indeed markedly increase from pyridine to pyrimidine. These experimental results are compared with state-of-the-art core-level photoionization calculations, which reveal the presence of several shape resonances in this low kinetic-energy region. The comparison of experiment and theory reveals that the photoelectrons originating from the nitrogen K-shell are trapped with increasing efficiency in the molecular plane in the sequence pyridine, pyrimidine, -triazine, i.e., with increasing number of nitrogen atoms. Additionally, the increasing symmetry of the molecular scaffold (from C2v in pyridine to D3h in -triazine) reduces the coupling among the photoelectron partial waves, which frustrates the decay of the trapped partial waves with high angular momentum. Pyrimidine is a building block of the nucleobases thymine, cytosine and uracil. All three azabenzene molecules are widespread structural motifs in biomolecules, drugs and molecular optoelectronics [21, 22, 23, 24]. The discovered mechanisms therefore suggest guidelines for extending attosecond chronoscopy to biomolecules and molecules used in optoelectronics.
II Experimental setup and data analysis
@@
Figure 1 illustrates the principle of the present experiments, carried out at the time-resolved atomic, molecular, and optical science (TMO) experimental hutch [25] of the Linac Coherent Light Source (LCLS). Linearly polarized attosecond X-ray pulses with a tunable central energy were obtained using the XLEAP technique [17] and were superimposed with circularly polarized short-wave infrared (IR, 1.3 m, 40 fs) pulses in the interaction region of a co-axial velocity-map-imaging (c-VMI) spectrometer. A continuous molecular beam containing pyridine, pyrimidine or -triazine intersected with the overlapping X-ray and IR pulses in the interaction region. The photoelectrons created from nitrogen K-shell ionization of the sample were streaked by the IR pulses and projected onto a position-sensitive imaging detector with a central hole. A photoelectron image recorded by ionizing pyrimidine with X-ray pulses centered at 454 eV is shown in Fig. 1A, whereby the inner (outer) ring corresponds to ionization of the nitrogen (carbon) K-shell electrons. Figure 1C shows how the relative photoionization delays between the N-1s and C-1s photoelectrons were determined. Because each photoelectron was angularly streaked into the direction of the IR-laser vector potential at the instant of photoionization (Fig.1B), the relative streaking angle between N-1s and C-1s electrons directly encodes their relative photoionization delay. The conversion from relative streaking angle to the relative photoionization delay is simply given by , where is the period of the streaking laser [26, 27, 28, 29, 30, 31, 32].
@@
Figure 2 illustrates how the relative streaking angles were obtained from the experimental data. Due to the large temporal jitter (500 fs) between the arrival time of XFEL and IR laser pulses [33], the streaking directions of the core-shell photoelectrons were completely random from shot to shot. The information of relative photoionization time delay is therefore only encoded in the relative streaking angle of the N-1s and C-1s electrons, which is recorded on a shot-by-shot basis. Figure 2A shows the difference image between the average over all shots corresponding to a narrow range of streaking directions sorted by post-selection of the images and the average over all streaked images without any sorting and selection. An angular shift between the streaking directions of C-1s (outer) and N-1s (inner) electrons can be observed, which directly reflects the relative photoionization delays, as described above. To accurately determine the relative streaking angle, we employ a partial covariance analysis that utilizes the large shot-to-shot variation of streaking directions to directly extract without estimating and sorting the streaking directions of C-1s and N-1s electrons in every single image. For a given relative streaking angle , the streaking-induced signal fluctuations in C-1s electrons at have the most positive covariance with those in N-1s electrons around . By correctly calculating the partial covariance coefficients between streaking-induced signal changes in N-1s and C-1s electrons, the differential streaking angle manifests itself as a global shift in the partial covariance maps, as shown in Fig. 2C. Additional details on the partial covariance analysis are given in Section 2.2.1 of the supplemental material (SM) and systematically explained in Ref. [34]. By comparing Figs. 3A,C to B,D, a relative photoionization delay of up to as is observed at eV, as opposed to a negligible delay at eV. We have also developed a complementary data-analysis method (the so-called “center-of-mass” method), which is described in Section 2.2.2 of the SM. These two independent analysis methods yielded consistent photoionization delays (Fig. S14), which confirms the reliability and robustness of both methods.
@@ III Theoretical calculations
Since such shape resonances are notoriously sensitive to electron correlations and molecular structure [36, 4, 9, 8], we have carefully benchmarked the core-level photoionization calculations against measured photoionization cross sections (Figs. 3D and S22), as well as photoelectron angular distributions measured in the present work (Figs. 3B and S23). Comparing the calculations using the equivalent-core (ECO) method with the experimental photoionization cross section (Fig. 3D) revealed very good agreement for the resonance lying below the nitrogen K-edge and the lowest-lying shape resonance after shifting the calculations to lower energies by 4.3 eV. In comparison, the correlation-polarization (CP) calculations only needed to be shifted by 0.6 eV to agree with the experimental cross sections and the calculated width of the lowest-lying shape resonance was in better agreement with the experiment than the ECO calculations. The CP calculations also agree better with the angular asymmetry parameters measured in this work (Figs. 3B and S23). Since both calculations overestimate the energetic position of the upper shape resonance by 3.5 eV, we therefore identified the CP calculations as the more accurate ones and from hereon exclusively discuss these results.
@@
Figures 3C,D demonstrate the two shape resonances of pyrimidine within the energy range, which lead to the local maxima of the cross section. The 1A′ symmetry component is dominating in the cross section, which displays an anti-bonding character and corresponds to the electron trapping that occurs in the molecular plane in both resonances. The insets in Fig. 3C show the one-electron representations of the corresponding resonant wavefunctions (Kapur-Peierls states, see SM for definition). The lower-lying shape resonance features 5 nodal planes and the upper one 6, showing that partial waves of high angular momentum are dominating the photoelectron trapping in both cases. The narrower width of the lower resonance nicely correlates with a longer photoemission delay and the broader linewidth of the upper one correlates with the shorter delay. The partial cross section corresponding to a total final-state wavefunction of 1A′′ symmetry, i.e., anti-symmetric with respect to the molecular plane and thus a -type continuum wavefunction, does not feature any sharp shape resonances above the ionization threshold (404 eV) in the CP calculation. Note however, the shifted ECO calculation does have a narrow feature slightly below the threshold, as shown in Fig. 3D, which is in agreement with the autoionization feature seen in a previous experimental study [35]. Thus we conclude that in the continuum above the nitrogrn K-edge the photoelectron trapping is dominated by the -type resonances in the molecular plane of pyrimidine. The probability density of the resonant wavefunction of -triazine at 418 eV is illustrated in the inset of Fig. 3E and in Fig. 4. Similar resonances occur in all three molecules, as shown in Fig. S20.
@@
Since our measurements extend to within 2 eV of the N-1s ionization threshold, they are particularly sensitive to the CLC delays. It has previously been shown that CLC delays in attosecond streaking experiments have the same origin of the continuum-continuum (CC) delays known from reconstruction of attosecond beating by interference of two-photon transitions (RABBIT) measurements [3]. The most commonly used expressions for the CC delays are reported in Ref. [37], which is based on the hydrogen-like atoms, and the radial wavefunctions are approximated considering the asymptotic phase shift (denoted as “P”) or with additional consideration of the long-range amplitude correction (denoted as “P+A”). The approximation of the hydrogenic radial wavefunction causes errors compared to the time-dependent Schrödinger equation (TDSE), particularly in the low-kinetic-energy region [37], which is relevant in our work. Recently, Ji et al. [38] reported using the analytical formula with hydrogenic radial wavefunctions (confluent hypergeometric functions) [39] instead of their asymptotic forms to calculate the CC transition of the hydrogen-like atoms, which shows excellent agreement with the TDSE in the whole kinetic-energy range. Here, we have adapted the same approach for calculating the CLC delay with a slight modification, where for the CC transition the increase or decrease of the kinetic energy is a continuous quantity instead of being equal to the energy of one IR photon, and the time delay is defined based on its derivative. The details can be found in Section 4.2 of the SM. This new approach yields very good agreement to the TDSE values reported in Ref. [31] (see Fig. S24A). Therefore, we used those results in Figs. 3 and 5. We also investigated the role of the Auger-Meitner decay and the post-collision interaction (PCI) between the Auger-Meitner electron and the photoelectron. We find that the correction of the Coulomb delay from a singly- to a doubly-charged cation is nearly compensated by the delay that is caused by the PCI effect (Fig. S25). Since these corrections are small and do not affect the relative delays between the molecular species that we studied, we refer the interested reader to the SM, and neglect these corrections for the remainder of the present discussion.
IV Discussion
@@
These results open the novel opportunity to analyze the mechanisms that govern the trapping of the photoelectron on the attosecond time scale. For this purpose, we analyze the properties of the trapped photoelectron wave function and the potential that traps it. Figure 4 compares the trapping of the outgoing photoelectron wave in the resonances at around 410 eV of pyridine and -triazine. The atomic cores provide attractive potentials in their vicinities, while the centrifugal energy creates a repulsive region centered on the aromatic ring. Their addition results in a potential barrier (black curves in Fig. 4) which traps the resonant wavefunction. The calculations provide a lifetime of 150 as for pyridine and 360 as for -triazine at this shape resonance. The stark difference in lifetime is reflected in the time-delay difference in the same energy region. The photoelectron wave initially originates from the 1s shells of the nitrogen atoms and then scatters in the molecular potential, which results in comparable amplitudes of the resonant wave function on all atoms of the molecular ring. But what causes the increase of the photoionization delays and resonance lifetimes with increasing number of nitrogen atoms and increasing symmetry?
@@
To answer this question, Figure 5A shows a sketch of effective radial potentials which indicate the height of the angular-momentum barrier for different values of the angular momentum, , and Figs. 5C,E,G show the absolute squares of resonant one-electron wavefunctions, resolved into partial waves of angular momentum and irreducible representations of the respective point groups (specified in each panel). Figure 5B shows the CLC delays and Figs. 5D,F,H show the calculated delays for all three molecules. We will focus the remainder of this discussion on the shape resonance located at about 410 eV in each molecule.
@@
In all three molecules, the partial-wave components with dominate in probability density, which is the consequence of the high barrier in the radial potential (panel A) arising from the centrifugal-potential contribution , where is the mass of the electron and its distance from the center of charge. The effective potentials of different partial waves are indicated by the shaded areas with different colors in the corresponding panels. It is clear that the main part of the wavefunction is confined in the quasi-bound region in the vicinity of the N and C atoms, while a tail tunnels through the barrier. In the case of pyridine (Fig. 5A), the dominant partial wave is least-well trapped compared to the other molecules, which is clear from comparing the modulations of the probability density at the barriers. This identifies tunneling through an angular-momentum barrier as the first mechanism governing electron trapping in the azabenzene series and shows that the trapping strength of the dominant partial wave increases in the order pyridine, pyrimidine, -triazine. This shows that substituting a C-H group with the more electronegative N atom increases the trapping strength and therefore the lifetime of the resonance. We note that the divergence of the Kapur-Peierls state for large is a consequence of their mathematical nature, specifically their complex-valued momentum, see SM Section 3.2.
Figure 5 also reveals a second mechanism at play. In addition to the changing probability density ratio between the wavefunctions in the quasi-bound and tunneling regions, we also observe the manifestation of molecular symmetry. The reason that the partial wave prevails in the resonant wavefunctions can be qualitatively understood from the inset panel in Fig. 5H. The resonant wavefunction displays five nodal surfaces, as qualitatively illustrated in the inset of Fig. 5H, corresponding to a dominant partial-wave character. However, the wavefunction also contains contributions from the other partial waves. As the symmetry of the molecular scaffold of the neutral molecule decreases from -triazine (D3h) to pyrimidine and pyridine (C2v), the relative amplitudes of to partial waves increases as a consequence of lower symmetry. The increasing relative amplitudes of the untrapped partial waves cause a faster decay of the resonant state.